monkeymode365:

pls help me Container A is cylinder with a radius of 6 units and a height of 6 units. A right cone has been carved from its base and has a height of 6 units. Container B has the same radius as container A. Which statement derives the formula to find the volume of container A? container A is a right cylinder that has had a right cone subtracted from its base, container B is half of a sphere 1 over 3π(62)(6) − π(6^2)(6) 2[1 over 3π(62)(6) − π(6^2)(6)] 2[π(62)(6) − 1 over 3π(6^2)(6)] π(62)(6) − 1 over 3π(6^2)(6)

AZ:

so container A is a right cylinder that has a right cone subtracted from its base Can you screenshot the image for more clarity? otherwise, the volume of container A is going to be (volume of right cylinder) minus (volume of right cone) do you know the formula for volume of a cylinder and volume of a cone?

monkeymode365:

1 attachment
AZ:

Good, that image helps Do you know the formulas? \( \text{Volume of cylinder} = \pi r^2 h\) \( \text{Volume of cone} =\dfrac{1}{3} \pi r^2 h\) remember what I said about the volume for your question? It's volume of the cylinder minus volume of cone can you plug in r = 6 and h = 6 what is your final answer?